OpenStudy (anonymous):

the length of a rectangle is 2 less than times its 3 width. Find the length if the perimeter is 76cm

OpenStudy (anonymous):

3w = L-2 then w = (L-2)/3 and perimeter 2(w+l) and substitute to get the answer